OpenStudy (anonymous):

What linear speed must a 4.5×10^−2 kg hula hoop have if its total kinetic energy is to be 0.20J ? Assume the hoop rolls on the ground without slipping. Express your answer using two significant figures.

OpenStudy (anonymous):

I tried 1/2mv^2 = 0.20 v^2=0.20*2/m =0.40 * 100/ 4.5 = 8.89 v=sqrt 8.89 = 2.98 m/s But that was wrong.